# 0.1 Simple Estimation of Bi-Lipschitz Constants at Initialization

### v1

@@ 0.6.1 Two Moons
Table 2 provides details on the accuracy of the learned model with different $\alpha$ and $\beta$, used for the grid search. As for the accuracy, the value of $\alpha$ does not matter as long as $\beta$ is chosen large enough. Indeed, a BLNN with $(\alpha,\beta)=(0.0,3.0)$ theoretically includes that with $(\alpha,\beta)=(1.0,2.0)$. That is why they exhibit similar performance. Nevertheless, a too loose setting of these parameters, especially for $\beta$, still seems to affect the training, and a as tight as possible choice is beneficial as we can see for $(\alpha,\beta)=(2.0,5.0)$. In a sense, decreasing $\beta$ which is mainly in charge of the Lipschitz constant helps better generalization as pointed out in prior work as well (B2017spectrally). In the same direction, a too large value of $\alpha$ leads to a nonsensical result as we can understand when $\alpha=5$. In this case, the model is randomly guessing which means that it is unsure everywhere. This is also translated in the uncertainty estimation as shown in Figure 37. The area where the model is unsure increases as $\alpha$ is increased, meaning that it becomes more sensitive to unknown data (see Figure 41). Comparatively, setting $\alpha=0$ does not affect so much since the injectivity constraint is already beneficial in this type of task. Finally, a too low value of $\beta$ means that the inverse Lipschitz and Lipschitz constants are close to each other and the function behaves more like a linear function. Consequently, learning becomes difficult as we can conclude from the table with $\beta=0.1$.

### v2

@@ 0.6.1 Two Moons
Table 2 provides details on the accuracy of the learned model with different $\alpha$ and $\beta$, used for the grid search. As for the accuracy, the value of $\alpha$ does not matter as long as $\beta$ is chosen large enough. Indeed, a BLNN with $(\alpha,\beta)=(0.0,3.0)$ theoretically includes that with $(\alpha,\beta)=(1.0,2.0)$. That is why they exhibit similar performance. Nevertheless, a too loose setting of these parameters, especially for $\beta$, still seems to affect the training, and a as tight as possible choice is beneficial as we can see for $(\alpha,\beta)=(2.0,5.0)$. In a sense, decreasing $\beta$ which is mainly in charge of the Lipschitz constant helps better generalization as pointed out in prior work as well (B2017spectrally). In the same direction, a too large value of $\alpha$ leads to a nonsensical result as we can understand when $\alpha=5$. In this case, the model is randomly guessing which means that it is unsure everywhere. This is also translated in the uncertainty estimation as shown in Figure 37. The area where the model is unsure increases as $\alpha$ is increased, meaning that it becomes more sensitive to unknown data (see Figure 41). Comparatively, setting $\alpha=0$ does not affect so much since the injectivity constraint is already beneficial in this type of task. Finally, a too low value of $\beta$ means that the inverse Lipschitz and Lipschitz constants are close to each other and the function behaves more like a linear function. Consequently, learning becomes difficult as we can conclude from the table with $\beta=0.1$. Similar trends could be also observed in more complex tasks such as FashionMNIST vs MNIST and FashionMNIST vs NotMNIST.
@@ 0.6.2 FashionMNIST
0.7 Partially Monotone Settings
Table 5 provides the experimental results with full data. We also compare with the certified monotone neural network of L2020certified (Certified) and the constrained monotone neural network of R2023constrained (Constrained). See Tables 2 and 3 of K2023scalable for a comprehensive comparison with other general models.
Table 5: Comparison of our model with state-of-the-art monotone models. Means and standard deviations over three trials. Results of prior models are from the original papers. C = COMPAS, BF = BlogFeedBack, LD = LoanDefaulter, HD = HeartDisease, AM = AutoMPG, Acc. = accuracy.
[…]
Constrained 69.2 $\pm$ 0.2 0.154 $\pm$ 0.001 65.3 $\pm$ 0.0 89.0 $\pm$ 0.0 8.37 $\pm$ 0.1
LMN 69.3 $\pm$ 0.1 0.160 $\pm$ 0.001 65.4 $\pm$ 0.0 89.6 $\pm$ 1.9 7.58 $\pm$ 1.2
SMNN 69.3 $\pm$ 0.9 0.150 $\pm$ 0.001 65.0 $\pm$ 0.1 88.0 $\pm$ 4.0 7.44 $\pm$ 1.2
Ours 69.4 $\pm$ 0.6 0.157 $\pm$0.001 65.5 $\pm$ 0.1 90.2 $\pm$ 1.6 7.13 $\pm$ 1.2
[…]